Journalist Report Jan 11

Crew 219 Journalist Report 11 JAN 2020

Author: Alejandro Perez, Crew 219 Engineer

Sol 6

Taking a Moment

Sol 6 has been one of the most rewarding days this mission. The day started with a beeping radio that woke me from my slumber. However, this was fine as we had a lot of work to do throughout the day. First was checking readings around the Hab and testing the new heating

system for the loft water tank pump, which worked perfectly. I conducted rounds shortly after the first EVA team was out in the field. The sun made a front stage appearance in today’s activities; providing warmth to the Greenhab and the surrounding desert. Later in the day on my EVA, this source of energy loosened the dirt beneath my feet making it difficult to walk around where there was no snow. Despite the continuous shine of the sun, the temperature remained below freezing for most of the day.

The middle of the day was highlighted by Hannah’s (LSO) cooking of homemade cornbread and chili. The cornbread was as soft as freshly fallen snow and the chili as tasty as mom’s home cooking. This was a welcome preparation for the second EVA team before going out in the field. The second EVA was set to explore Tank Wash and the valley area near Murphy’s Canyon. Throughout the expedition, Abdul (GEO) and I found crystalline rocks that glistened in the sunlight catching our eyes like a minnow to a barracuda. Next to these clear rocks were tracks of what looked like an intense chase between a mountain lion and deer. At this time in the EVA, we were close to our sampling location where we found possible sources of lichen on the rock bed above the valley. Finally, we have found possible extraterrestrial life on this Red Planet. The team then climbed a short way to collect

additional samples higher up and the views were breathtaking. As far as the eye can see were rivers of snow and sediment washing through the canyon like a lost land locked in time. Each mountain seemed to be a smooth carving of what looked like giant pieces of chocolate and mousse cake. We took a couple minutes to enjoy the moment and ended the hike with awkward EVA suit hugs and laughs at the top of this climb.

The euphoria of this journey had the EVA team in a nice gallop on the way back to the rovers. At this point in the day, the sun had done its job of waking up the land surrounding us. Unfortunately, the camera the team had brought on the EVA had died just before the sampling site. I guess this will be a moment shared only with people standing next to me. For someone from the concrete jungle of the Northeast, this has to be the most rewarding and life-changing experiences I have ever been a part of.

Science Report Jan 11

Science Report 11-Jan-2020

Crew Science Officer: Hannah Blackburn

1. Biometrics and Neurobehavioral Research

Continued taking sleep and post-EVA logs. Post-EVA biometrics were recorded.

2. UAV

Nothing to report.

3. Dust Mitigation for Optical Mirrors

Mirrors were examined in the Science Dome. Samples were selected for simulated dust storm and dried in the oven.

4. Astrophotography of Celestial Bodies

3 observations were taken last night. We will download the results between 2 and 6 am tomorrow.

5. Remediation of Mars Regolith

Continued examining regolith under the microscope. One dead algae was found, but no living organisms.

6. Chemical and Mineralogical Composition of the MDRS Site

Samples were dried in the oven using glass beakers. Additional samples were taken during EVAs 11 and 12.

7. Protocols for the Discovery of Life on Mars

Additional samples collected during EVA 12.

Glassware check out:

10 glass beakers (sizes between 200 and 500 ml) were used to dry regolith samples.

A reminder to all crewmembers: There’s a $300 fine for using any glassware material without MDRS Mission support permission.

Green Hab Report

Crew 219 GreenHab Report 11-JAN-20

Crew GreenHab Officer: Cynthia Montanez

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 25.2 °C; 17.3 %

Average Max: 28 °C, 20.3%

Average Min: 19.8 °C, 16%

11:14 A.M.

Floor Unit: 20 °C

Electronic: 27.8 °C

humidity 16%

Max: 28.8 °C; 21%

Min: 17.8 °C; 16%

2:51 P.M.

Floor Unit: 14 °C

Electronic: 24.8 °C

humidity 18%

Max: 27.5 °C; 20%

Min: 20.9 °C; 16%

5:38 P.M.

Floor Unit: 14 °C

Electronic: 23.1 °C

humidity 18%

Max: 27.7 °C; 20%

Min: 20.7 °C; 16%

Hours of supplemental light: Light system 7:00PM-12:00AM.

Daily water usage for crops: 9.20 Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – ~179.45 Gal.

Time(s) of watering for crops:

11:14 A.M.

SEASONAL PLANTING HISTORY:

Change to crops: Acorn pumpkins are flowering and more sprouts are appearing on the cherry tomatoes.

OTHER NOTES:

1st sprouts: N/A

Harvest: N/A

Other: Today At 11:14 A.M. I fed the plants. Also, at 2:51 P.M., I saw that the spray helped in removing the gnats from the icicle radishes, but they made their way to the purple carrots. I sprayed the carrots with the same spray in attempts to kill the gnats. At 5:38 P.M. I went to check on the carrots and I saw no more gnats.

Operations Report Jan 11

Crew 219 Operations Report 11-JAN-2020

Sol: 6

Name of person filing report: Alejandro Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 14.5

From what time last night: 17:00

To what time this morning: 07:30

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 77%

08:40 = 96%; 09:30 = 94%; 10:20 = 92%; 11:00 = 90%; 11:30= 89%; 12:00=

89%; 12:30= 88%; 13:00 = 85%; 13:30= 85%; 15:00= 82%; 16:00= 80%;

17:00= 77%

Diesel Reading – 65%

Station Propane Reading- 45%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0147567,5

Water (static tank) (gal) – 444

Static to loft Pump used – Yes

Water in Green Hab (gal): 179.45

Water in Science Dome (gal): 0

Toilet Tank Emptied – Yes

Diemos rover used: Still in the Work shop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- Yes

Starting Hours: 124.5, 124.5

Beginning Charge: 100%, 100%

Ending Hours: 124.5, 124.9

Ending Charge: 23%, 47%

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

Starting Hours: 129.9

Beginning Charge: 100%

Ending Hours: 130.3

Ending Charge: 39%

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where?

CrewCar used and why, where?

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: Suit #5 was tested today and worked properly.

Summary of Hab: Air filter for Heater has been replaced with a Clean one. New Teflon tape was used on the Loft tank feedline.

Summary of Science Dome operations: Temperatures are hovering around 7°C during the day.

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: Metal dish scrubber and a brush to wash dishes for pots and pans.

EVA Report Jan 11

Crew 215 EVA Report 15-Nov-2019

EVA #11

Author: Cynthia Montanez

Purpose of EVA: Geologic Sample Collection

Start Time: 9:50 A.M.

End Time: 10:40 A.M.

Narrative: EVA #11 was led by Cynthia Montanez (GHO). Her crew’s destination was south of Marble Ritual for a quick geologic sample collection. Upon reaching the destination, the GHO identified 5 sample sites that were 10 feet apart. Each sample site had its coordinates

recorded on the GPS. Once completed, the crew made their way back to the Hab for the 5-minute depressurization at 10:35 A.M.

Destination: South of Marble Ritual

Coordinates: 0518600/4250300

Participants: Cynthia Montanez (GHO), Hannah Blackburn (LSO), Keith Crisman (HSO), Dave Masaitis (CO)

Roads and Routes: Directly east out of Hab past Cow Dung Road

Mode of Travel: Walking

Crew 219 EVA Report 11 JAN 2020

EVA#: 12

Author: Nathan Hadland

Purpose of EVA: The primary purpose of this EVA was sampling for our geology project. Post-EVA, we collected data for our biometrics project.

Start time: 1335 hrs

End time: 1553 hrs

Narrative: Due to the assistant director using the rovers, the EVA was delayed by 30 minutes to allow Spirit to achieve 100% SOC. At 1335 hours, EVA #12 initiated a 5-minute decompression and moved to Curiosity and Spirit. After reporting SOC and hours, we progressed north along Cow Dung Road until we reached Tank Wash. We moved east on

foot through Tank Wash until just above Murphy’s Canyon. After taking 5 regolith samples and 3 biological samples (lichen), the crew then then turned around and returned to the rovers. Curiosity and Spirit proceeded safely back to the Hab at 1548 hours and the EVA crew successfully pressurized in the air lock at 1553 hours.

Destination: Tank Wash

Coordinates: 0519400/4253800

Participants: Nathan Hadland (XO), Alejandro Perez (ENG), Robinson Raphael (ASTRO), Abdul Elnagdi (GEO)

Roads and routes per MDRS Map: North along Cow Dung Road until 518100/4253550 and then proceeding East on foot through Tank Wash until just above Murphy’s Canyon.

Astronomy Report Jan 11

Crew 219 Astronomy Report 11 Jan 2020

Crew Astronomer: Robinson Raphael

Crew: 219

Date: Jan 11th, 2020

MDRS ROBOTIC OBSERVATORY: (MDRS-14)

MDRS-14 took observations of:

NGC 1068: (4 exposures of 120 seconds in B, 4 exposures of 60 seconds in V, 4 exposures of 30 seconds in R)

NGC 7318: (4 exposures of 120 seconds in B, 4 exposures of 60 seconds in V, 4 exposures of 30 seconds in R)

M32: (4 exposures of 120 seconds in B, 4 exposures of 60 seconds in V, 4 exposures of 30 seconds in R)

Already have submitted observations for NGC 2623, NGC 2960, NGC 4258, Owl Nebula, Whirlpool Galaxy, and Crab Nebula

Images submitted with this report: None (Still have to calibrate and stack the images with AstroImageJ)

Problems Encountered: Flats for the each of the filters (B,V,R) are not recorded and still experimenting with the exposure times for certain celestial bodies.

Sol Summary Report Jan 11

Crew 219 Sol Summary Report 11-01-2020

Sol: 6

Summary Title: A Clear Day of Great Adventure

Author’s name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: For the first time since touching down on Mars, Crew 219 awoke to clear skies and a visibly rising sun. Thanks to those clear skies, our R. Raphael actually got astronomical
observations! EVA #11 moved out first, making quick work of navigation and sample collection. Samples were transferred to the Science Dome for analysis, and after a lunch of fresh cornbread and assorted leftovers, EVA #12 took Spirit and Curiosity north to yet unexplored terrain. After their departure, more sample analysis continued in the science dome, and to everyone’s joy EVA #12 returned with some samples of yet-unidentified potential life forms! For a crew from the
Astrobiological Research and Education Society, potential biological samples bring a whole new form of excitement. H. Blackburn is making gnocchi for dinner, and the whole crew is looking forward to a
Command-mandated DONSA!!! (Day Of No Scheduled Activities)

Look Ahead Plan: DONSA – rest, recreation, and perhaps some maintenance if necessary

Anomalies in work: None

Weather: Cold and clear, with daytime temperatures hovering around freezing

Crew Physical Status: Nominal

EVA:

EVA #11 – S. of Marble Ritual (0518600/4250300)

EVA #12 – E of Tank Wash (0519400/4253800)

Reports to be filed:

Sol Summary
EVA Report 11 & 12
Astronomy Report
Operations Report
GreenHab Report
Research Report
Journalist’s Report
Photos (Including Photo of the Day)

Support Requested:
– Refer to Operations Report